At the present time, Jenifer is 20 years younger than her mother. In 4 years, the ratio of Jenifers age will be 4 to 9. Find Jenifers present age
Answer by checkley71(8403)

X-Y=20 OR X=Y+20 THEN (Y+4)/(Y+20+4)=4/9OR 9(Y+4)=4(24+Y) OR 9Y+36=96+4Y OR
5Y=60 OR Y=60/12 OR Y=12 IS THE AGE OF JENIFER
PROOF (12+4)/(12+20+4)=4/9 OR 16/36=4/9 OR 144=144